# A Simple Apparatus for Testing PMT Humidity Tolerance

### v1

@@ 1 Introduction
In order to test the operability of this PMT model in high humidity, we constructed a low-cost test setup featuring a watertight box into which we injected high-humidity air, and which allowed slight elevation of temperature. Section 2 describes this environmental chamber in detail. In Sec. 4 we show an example of test data acquired with this chamber, and summarize its performance.
@@ 2.1 Humidity injection
Figure 4: A typical time series of humidity measurements with no PMT within the plastic container, taken by the humidity/temperature sensor in the plastic container and recorded by the Arduino. The Vicks humidifier is turned on at time $t=0$. A condensation blockage in the PVC tube was cleared at about minute 55. The error bars show the manufacturer-specified relative accuracy of ±$\pm 2\%$. For orientation, the typical PMT run is one minute long.
@@ 2.2 Operation at elevated temperature
Figure 5: A typical set of temperature measurements with no PMT within the plastic container, taken by the humidity/temperature sensor in the plastic container and recorded by the Arduino. The ceramic heat emitter is turned on at time $t=0$. Although the digital temperature controller (in the dark box) is set to 30 C, the humidity/temperature sensor in the nested plastic box typically reads slightly lower temperatures as shown. The error bars show the manufacturer-specified accuracy of ±$\pm 0.2^{\circ}$.
@@ 3 PMT Testing in Environmental Chamber
The PMT was also operated at a slightly elevated temperature. Once a stable temperature of about 26.5C was reached inside the plastic box under conditions of elevated humidity, the PMT’s response to LED pulses was observed over the course of 5 days. No statistically significant change in the PMT response was observed.

### v2

@@ 1 Introduction
In order to test the operability of this PMT model in high humidity, we constructed a low-cost test setup featuring a watertight box into which we injected high-humidity air, and which allowed slight elevation of temperature. Section 2 describes this environmental chamber in detail, while Sec. 3 shows an example of test data acquired with the chamber. In Sec. 4 we summarize cost and performance.
@@ 2.1 Humidity injection
During very long runs, e.g. the 5-day run described in Sec. 3, condensation buildup becomes both less frequent and less predictable compared to the typical startup period shown in Fig. 4. With checks of the humidity level a few times a day and interventions as needed, the relative humidity level could be kept above about 70%.
Figure 4: A typical time series of humidity measurements with no PMT within the plastic container, taken by the humidity/temperature sensor in the plastic container and recorded by the Arduino. The Vicks humidifier is turned on at time $t=0$. A condensation blockage in the PVC tube was cleared at about minute 55. The error bars show the manufacturer-specified relative accuracy of $\pm 2\%$. For orientation, the typical PMT run is one minute long.
@@ 2.2 Operation at elevated temperature
Figure 5: A typical set of temperature measurements with no PMT within the plastic container, taken by the humidity/temperature sensor in the plastic container and recorded by the Arduino. The ceramic heat emitter is turned on at time $t=0$. Although the digital temperature controller (in the dark box) is set to 30 C, the humidity/temperature sensor in the nested plastic box typically reads slightly lower temperatures as shown. The error bars show the manufacturer-specified accuracy of $\pm 0.2^{\circ}$C.
@@ 3 PMT Testing in Environmental Chamber
The PMT was also operated at a slightly elevated temperature. Once a stable temperature of about 26.5C was reached inside the plastic box under conditions of elevated humidity, the PMT’s response to LED pulses was observed over the course of 5 days as a proof-of-concept measurement for long-term tests as described in Sec. 2.2. No statistically significant change in the PMT response was observed.